3.73(98.35x+47)= -96(70x -35.5)

it would help so much if you could answer this question for me.

Answer by stanbon(75887)

3.73(98.35x+47) = -96(70x -35.5)
----
366.8455x + 175.31 = -6730x -3408
7096.8455x = -3583.31
---
x = -0.5049
================
Cheers,
Stan H.